one leg of the right traingle is 5 iches longer than the other leg. if the hypotenuse measures 25 inches, find the length of each leg. Could you please show work. Thankyou
Answer by Cromlix(4381)

Hi there,
Make the 'other leg' = 'x'
So, the 'one leg' = x + 5
Now by Pythagoras the sum of the square of the
hypotenuse = the sum of the squares of the other
two sides.
x^2 + (x + 5)^2 = 25^2
Multiply out
x^2 + x^2 + 10x + 25 = 625
Collect like terms
2x^2 + 10x + 25 - 625 = 0
2x^2 + 10x - 600 = 0
Divide by 2 through out
x^2 + 5x - 300 = 0
Factorise
(x - 15)(x + 20) = 0
x + 20 = 0
x = -20 (no answer as -ve)
...........
x - 15 = 0
x = 15
Size of legs:
'Other leg' = 15 inches
the 'One leg' = 20 inches
Hypotenuse = 25 inches
15^2 + 20^2 = 25^2
225 + 400 = 625
625 = 625
Hope this helps:-)